An airplane is flying in air with a density of 1.29 kg/m3 . A pressure gauge measures the difference in pressure between a point on the nose of the airplane and a point on the fuselage. The pressure on the nose is higher by 60.0 kPa. What is the speed of the plane relative to the air?

Please explain the process! Thanks!

A) 295 m/s B) 300 m/s C) 305 m/s D) 310 m/s E) 315 m/s
Answer #1

C) 305 m/s

here is the explanation.


Apply Bernoulli's equation

P1 + 0.5*rho*v1^2 = P2 + 0.5*rho*V2^2

here, v2 = 0

so,

0.5*rho*v1^2 = P2 - P1

v1 = sqrt(2*(P2-P1)/rho)

= sqrt(2*60*10^3/1.29)

= 305 m/s